# Teaching Transformers Causal Reasoning through Axiomatic Training

### v2

@@ Abstract
For text-based AI systems to interact in the real world, causal reasoning is an essential skill. Since active interventions are costly, we study to what extent a system can learn causal reasoning from symbolic demonstrations of causal axioms. Specifically, we present an axiomatic training method where the system learns from multiple demonstrations of a causal axiom (or rule), rather than incorporating the axiom as an inductive bias or inferring it from data values. A key question is whether the system would learn to generalize from the axiom demonstrations to more complex scenarios. Our results, based on applying axiomatic training to learn the transitivity axiom and d-separation rule, indicate that such generalization is possible. To avoid data contamination issues, we start with a 67 million parameter transformer model and train it from scratch. On both tasks, we find that a model trained on linear causal chains (along with some noisy variations) can generalize well to complex graphs, including longer causal chains, causal chains with reversed order, and graphs with branching. To handle diverse text inputs, the same method is extended to finetune language models. Finetuning Llama-3.1 8B model on our axiomatic data leads to significant gains on causal benchmarks such as Corr2Cause and CLEAR, in some cases providing state-of-the-art performance surpassing GPT-4.
@@ 1 Introduction
Causal reasoning can be defined as a set of reasoning procedures consistent with pre-defined axioms or rules that are specific to causality (Galles & Pearl, 1997). For instance, under stable causal models, the transitivity axiom (“if A causes B and B causes C, then A causes C”) helps answer questions of cause and effect between pairs of variables in a system. Similarly, the d-separation rule connects independence of variables and their causal graph structure, and forms the basis of many graph discovery and effect identification algorithms. Given a causal task and data observations from a system, axioms or rules are typically incorporated as inductive biases in a machine learning (ML) algorithm, through regularization, model architecture, or the choice of variables for a particular analysis. Depending on the kind of available data—observational, interventional, or counterfactual—Pearl’s ladder of causation (Bareinboim et al., 2022) defines the kinds of causal reasoning that is possible.
As axioms are the building blocks of causality, we study whether it is possible to directly learn the axioms or rules using ML models. That is, rather than learning from data that is the result of axioms obeyed by a data-generating process, what if a model can learn an axiom (and thus causal reasoning) directly from symbolic demonstrations of the axiom? This question gains relevance as language models make it possible to learn over symbolic data expressed in natural language. In fact, recent studies have evaluated causal reasoning capabilities of large language models (LLMs) by encoding causal reasoning problems in natural language (Kıcıman et al., 2023; Jin et al., 2024a, b). Our goal is to study whether directly teaching the axioms can be a viable way to improve causal reasoning of language models.
@@
To answer this question, we consider a setup where a model is trained on axiomatic demonstrations over simple chain-like graphs of size 3-6 nodes and evaluated on more complex graphs, including longer chain-like graphs of size 7-15, graphs with branching, longer variable names, and edge direction perturbations. To avoid any contamination concerns with the pre-training data of an existing language model, we first train a transformer model from scratch. For both transitivity and d-separation, we find that a model trained on axiomatic demonstrations learns to apply the axiom multiple times to answer questions over more complex graphs. Diversity in the training data matters. For transitivity, a model trained only on simple directed chains generalizes to longer length chains, but is unable to generalize to graphs with branching or edge direction perturbations. In comparison, a model trained on a combined dataset of simple chains and chains with some edges randomly reversed, generalizes well across all kinds of evaluation scenarios. In particular, for d-separation, our 67 million parameter model outperforms billion-scale models such as GPT-4 under both zero-shot and multi-shot settings. Extending the findings on positional embedding for length generalization in NLP tasks (Kazemnejad et al., 2023; Bhattamishra et al., 2020; Haviv et al., 2022), we find that rotary position embedding works the best for causal generalization.
@@
Next, we study whether the same axiomatic training dataset can also help to improve causal reasoning of pre-trained large language models. We fine-tune Llama-3.1-8B-Instruct model over axiomatic datasets for transitivity and d-separation. We evaluate on two benchmarks: CLEAR (Chen et al., 2024b), that includes a test set for measuring d-separation capabilities; and Corr2Cause (Jin et al., 2024b) on inferring causal structure from correlational statements. Note that our model is not trained on any of these datasets. We find significant gains due to axiomatic fine-tuning: in CLEAR, the accuracy on d-separation increases from 30 to 70 % on Yes/No task, and goes from 33 to 50% on Multi-Choice Questions. On Corr2Cause, the F1 score improves significantly, increasing 2x compared to the baseline and outperforms GPT-4 (F1=0.29) by 8 p.p thus highlighting the impact of axiomatic training on complex real world datasets for causal reasoning.
@@
Our work provides a new paradigm of teaching models causal reasoning through symbolic demonstrations of axioms or rules, which we call axiomatic training. Such symbolic data can be cheaply generated for multiple axioms and added to the finetuning data of language models. More generally, our results contribute to the literature on causal learning from passive data (Lampinen et al., 2023), showing a general way to learn causal reasoning with passive demonstrations. Our code repository can be accessed at: https://github.com/AniketVashishtha/CausalAxioms
@@ 2 Related Work
LLMs for Knowledge-Driven Causal Reasoning: Recent developments in Large Language Models (LLMs) have highlighted their potential for knowledge-driven causal discovery. Unlike traditional methods which focus on statistical patterns or correlations, LLMs utilize knowledge acquired through their pretraining to reason about and identify causal structures based on metadata of variables (Kıcıman et al., 2023; Ban et al., 2023; Long et al., 2023; Willig et al., 2022; Vashishtha et al., 2023). However, possibility of memorization of existing benchmarks in the pretraining of these LLMs has been a major criticism. As a result, recent work (Zečević et al., 2023) argues that LLMs are not actually performing causal reasoning, but simply learning correlations about causal facts. In addition, there are critical failure modes of using LLMs for causal discovery due to hallucinations or not obeying the acyclic constraint when generating graph edges (Vashishtha et al., 2023). To evaluate causal reasoning capabilities of LLMs, (Jin et al., 2024b) and (Jin et al., 2024a) propose formal causal inference evaluation benchmarks to infer direct and indirect causal relationships, and highlight the failure of LLMs in performing accurate causal reasoning.
Impact of Positional Encoding on Generalization: Length generalization capabilities of transformers has been studied in the past to better understand their different failure modes across various settings (Hupkes et al., 2020; Zhang et al., 2023; Furrer et al., 2021). Previous work (Kazemnejad et al., 2023; Bhattamishra et al., 2020; Haviv et al., 2022; Shen et al., 2023) emphasizes the impact of positional encoding in length generalization capability of transformers. To understand how transformers can be optimized for learning through axiomatic training and generalizing to unseen larger causal structures, we also examine different types of positional encoding such as no positional encoding (PE), Learnable PEs (Radford et al., 2018) and Sinusoidal PEs (Vaswani et al., 2023).
@@ 3.1 Axioms of Causality: Transitivity
$(X\nrightarrow Y|Z)\implies{}$ $(A\nrightarrow Y|Z)$
${}\vee(X\nrightarrow A|Z)$
$\forall A\notin X\cup Z\cup Y$
which can be simplified using the contrapositive.
$\exists A\notin X\cup Y\cup Z$ $s.t.$ (1)
$\underbrace{(X\rightarrow A|Z)\wedge(A\rightarrow Y|Z)}_{P:\text{premise}}$
@@
Definition 3.3 (Definition 1.2.3 in Pearl (2009a)).
@@
Figure 1: Evaluating structural generalization of transformers through axiomatic training. We train a transformer on two simple causal structures: chains and chains with random flipping of some edges. All training instances consist of 3-6 nodes. The trained model is evaluated on significantly more complex structures: bigger causal chains with >6 nodes, general branched networks with higher average in-degree and out-degree, complete reversals, longer sequences, shuffled natural language statements of sequences and longer node names
@@ 4.1 Training Data: Diversity is key
We train the model on data from simple causal graphs such as sequential chains with 3-6 nodes and evaluate its performance on more complex graphs 1. To enhance generalization, we introduce structured perturbations in the training data across three axes: node names, causal structure types, and the number of nodes in the causal graph.
@@
      Sequential: All causal edges are directed forward, thus forming a typical chain DAG, e.g. X $\rightarrow$ Y $\rightarrow$ Z.
@@ 4.2 Tokenization, Training Loss & Architecture
We train the transformer model from scratch to ensure that the model has not seen such axioms in the pertaining step and thus requires a true correct understanding of axioms to perform well. Later we also tested on a pre-trained model fine-tuned on our dataset.
@@
Positional Encoding. In addition to the training data and loss function, recent work (Kazemnejad et al., 2023) has shown that the choice of positional encoding is important for generalizing a transformer to longer or complex inputs. Therefore, we experiment with different positional encoding to understand their impact on generalization in causal tasks: learnable (LPE) (Radford et al., 2018), sinusoidal (SPE) (Vaswani et al., 2023), rotary (RoPE) position encodings (Su et al., 2023), and no positional encoding (NoPE) (Kazemnejad et al., 2023; Haviv et al., 2022). See Appendix E for details.
Finetuning. Apart from training a transformer from scratch, we also fine-tune a pre-trained language model (Llama-3.1-8b-Instruct (gra, 2024)) on our axiomatic training data.
4.3 Evaluation setup: Assessing Axiomatic Learning in Transformers
@@
    Node Name Shift: Testing the model’s performance on longer node names, from 1-3 characters used in the training set to 8-10 characters. This is motivated by (Jin et al., 2024b) who show how change in node names results in generalization failure on causal tasks for language models.
@@
    Branching Factor: One of the most complex evaluation setups, with DAGs containing multiple branches, colliders, and forks. Let the branching factor be defined as the ratio between a number of edges and a number of nodes. Thus, the branching factor for the training set is $\leq 1$. Then, we create a different evaluation set with multiple densely branched networks constructed using the Erdös-Rényi model, with different branching factors.
Benchmark evaluation. To test whether such simple axiomatic training is helpful in more complex scenarios, we evaluate our models on existing causal reasoning benchmarks, Corr2Cause (Jin et al., 2024b) and CLEAR (Chen et al., 2024b). Models trained from scratch on axiomatic instances with limited vocabulary and capability cannot be tested on diverse datasets (due to out-of-vocabulary issues). Therefore we use finetuned language models on these datasets.
@@ 5.1 Training and Evaluation Datasets
Training Datasets. The training data consists of sequential chains of lengths from [3,6]. In addition to sequential chains, random flipping of edges is done with 0.5 probability. See Appendix F for details on these hyperparameters. Our training data consists of 175k axiom demonstrations. We use three versions of training data to evaluate the impact of different data perturbations.
@@
    Training Setup 1 (TS1): This setup comprises of 73k examples where the underlying graphs has random flipping and 101k causal graphs where the underlying graphs has sequential chains. Since flipping is done randomly across all consecutive pairs of nodes in the given chain, some complete reversals are also formed. In this training set, around 12k completely reversed chains are present.
@@
    Training Setup 2 (TS2): This setup comprises more simple sequential chains (132k), while we decrease chains with random flipping (42k) to keep the overall size around 175k to see the effect of adding examples with complicated graphs on model’s performance.
Evaluation Datasets. In addition to the evaluation sets described earlier (length generalization, node name shift, order of chains, and branching), we add another evaluation set that is a combination of three shifts.
$\mathbf{MultiEval_{SLR}}$ (Shuffling + Random Flipping + Length Sequence): This setup involves evaluation on 3 levels of complexities together: shuffling of sentence for representing the sequences, each sequence having random flipping, and some sequences having longer length than sequences in training set (upto 9).
5.2 Implementation Details: Architecture and Training Procedure and Positional Encoding
We train a decoder-based 67 million parameter model based on GPT-2’s architecture. The model has 12 attention layers, 8 attention heads and 512 embedding dimensions. The model is trained from scratch on each of our training datasets. All models are trained for 100 epochs using the AdamW optimizer with 1e-4 learning rate. We use three variant of positional encoding when training the transformer: SPE: Sinusoidal Positional Encoding (PE), LPE: Learnable PE, RoPE: Rotary PE, NoPE: without any PE.
5.3 Baselines Using Existing LLMs
Given recent work on how LLMs can be leveraged for causal reasoning (Kıcıman et al., 2023; Vashishtha et al., 2023; Ban et al., 2023), we include language models such as GPT-4 (gpt-4-32k) (ope, 2024), Gemini (gemini-pro) (gem, 2024) and Phi-3 (Phi-3-mini-128k-instruct) (abd, 2024) as baselines. Note that each of these models is significantly larger than our model and known to perform well on reasoning tasks, with the smallest baseline model Phi-3 having 3.8 billion parameters (Li et al., 2023).
Zero Shot Setting To evaluate the baseline models, we follow a simple zero-shot prompting strategy. For each tuple, we provide the natural language expression of the causal graph (Premise) followed by the question (Hypothesis) and prompt the LM to answer it in either ‘Yes’ or ‘No’ (Label). Here is an example prompt: “EX causes T. T causes 9. 9 causes W. W causes 7. 7 causes M. M causes a. Does EX cause T? Answer in ‘Yes’ or ‘No’ only”. See Table LABEL:tab:prompt_table contains examples of prompts used.
Multi Shot Setting (In-context-learning). Since these LLMs might not have seen such tasks before, we include some examples along with their true label in the prompt and then we add the evaluation example in the end. This ensures LLMs can do in-context learning and thus a fair comparison against our model that is trained explicitly on the axiomatic dataset. We present few-shot instances from our training set that includes sequential causal chains, along with a few examples with random flipping of edges. Appendix B.1 contains the multishot prompt used for querying baseline LLMs.
5.4 Results: Generalization to Complex Causal Scenarios

### v3

@@ Abstract
For text-based AI systems to interact in the real world, causal reasoning is an essential skill. Since active interventions are costly, we study to what extent a system can learn causal reasoning from symbolic demonstrations of causal axioms. Specifically, we present an axiomatic training method where the system learns from multiple demonstrations of a causal axiom (or rule), rather than incorporating the axiom as an inductive bias or inferring it from data values. A key question is whether the system would learn to generalize from the axiom demonstrations to more complex scenarios. Our results, based on applying axiomatic training to learn the transitivity axiom and d-separation rule, indicate that such generalization is possible. To avoid data contamination issues, we start with a 67 million parameter transformer model and train it from scratch. On both tasks, we find that a model trained on linear causal chains (along with some noisy variations) can generalize well to complex graphs, including longer causal chains, causal chains with reversed order, and graphs with branching. To handle diverse text inputs, the same method is extended to finetune language models. Finetuning Llama-3-8B-Instruct model on our axiomatic data leads to significant gains on causal benchmarks such as Corr2Cause and CLEAR, in some cases providing state-of-the-art performance surpassing GPT-4.
@@ 1 Introduction
Causal reasoning can be defined as a set of reasoning procedures consistent with pre-defined axioms or rules that are specific to causality (Galles & Pearl, 1997). For instance, under stable causal models, the transitivity axiom (“if A causes B and B causes C, then A causes C”) helps answer questions of cause and effect between pairs of variables in a system. Similarly, the d-separation rule (Pearl, 2009) connects independence of variables and their causal graph structure, and forms the basis of many graph discovery and effect identification algorithms. Given a causal task and data observations from a system, axioms or rules are typically incorporated as inductive biases in a machine learning (ML) algorithm, through regularization, model architecture, or the choice of variables. Depending on the kind of available data—observational, interventional, or counterfactual—Pearl’s ladder of causation (Bareinboim et al., 2022) defines the kinds of causal reasoning that is possible.
As axioms are the building blocks of causality, we study whether it is possible to directly learn the axioms or rules using ML models. That is, rather than learning from a dataset sampled from a data-generating process that obeys causal axioms, what if a model can learn an axiom (and thus causal reasoning) directly from symbolic demonstrations of the axiom? This question gains relevance as language models make it possible to learn over symbolic data expressed in natural language. In fact, recent studies have evaluated causal reasoning capabilities of large language models (LLMs) by encoding causal reasoning problems in natural language (Kıcıman et al., 2023; Jin et al., 2023, 2024). We study whether directly teaching the axioms can be a viable way to improve causal reasoning of language models.
Figure 1: Axiomatic Training for imparting causal reasoning to language models. Given an axiom, we construct a training dataset comprising <premise, hypothesis, conclusion> triplets based on simple chain-like graphs of 3-6 nodes. A transformer model trained from scratch on such instances generalizes to much more complex graphs, including longer causal chains with >6 nodes, branched networks with higher average in-degree and out-degree, complete reversals, shuffled statements, and longer node names. Moreover, when an existing model such as Llama-3-8B-Instruct model is trained on the same dataset, it leads to significant (up to 20 percentage point ($p.$)) improvement in accuracy on causal reasoning benchmarks such as CLEAR and Corr2Cause.
@@
To answer this question, we consider a setup where a model is trained on axiomatic demonstrations over simple chain-like graphs with 3-6 nodes and evaluated on more complex graphs, including longer chain-like graphs with 7-15 nodes, graphs with branching, longer variable names, and edge direction perturbations (see Figure 1). To avoid any contamination concerns with the pre-training data of an existing language model, we first train a transformer model from scratch. For both transitivity and d-separation, we find that a model trained on axiomatic demonstrations learns to apply the axiom multiple times to answer questions over more complex graphs. In particular, diversity in the training data plays a crucial role in enabling this generalization. For transitivity, a model trained on a combined dataset of simple directed chains and chains with some edges randomly reversed, generalizes well across all kinds of evaluation scenarios. However, a model trained only on simple directed chains fails to generalize to graphs with branching or edge direction perturbations. Moreover, for d-separation, our 67 million parameter model outperforms billion-scale models such as GPT-4 under both zero-shot and multi-shot settings. Extending the findings on positional embedding for length generalization in NLP tasks (Kazemnejad et al., 2023; Bhattamishra et al., 2020; Haviv et al., 2022), we find that rotary positional encoding performs the best for causal generalization, closely followed by no positional encoding.
@@
Next, we study whether the same axiomatic training dataset can also help to improve causal reasoning of pre-trained large language models. We fine-tune Llama-3-8B-Instruct model over axiomatic datasets for transitivity and d-separation. We evaluate on two benchmarks of causal reasoning in natural language: CLEAR (Chen et al., 2024) and Corr2Cause (Jin et al., 2024). The CLEAR dataset tests language models’ understanding of causal graphs through a wide range of binary and multi-choice tasks, including d-separation, Markov equivalence, and backdoor adjustment. Even though we don’t finetune on their dataset or question type, we find significant performance gains on the d-separation task after axiomatic fine-tuning: accuracy increases from 30 to 70% on Yes/No task, and goes from 33 to 50% on Multi-Choice Questions.
Corr2Cause is another challenging dataset to assess a language model’s ability to infer causal relationships (e.g., ancestor, collider, or fork) given correlational statements expressed in natural language. Solving this task requires knowledge of multiple causal concepts including d-separation, transitivity, Markov property and the Markov equivalence class. Jin et al. (2024) show how frontier models such as GPT-4 also fail at this task. We show that axiomatic finetuning leads to a substantial F1 improvement of nearly 20 percentage points over the Llama-3-8B instruct baseline and even surpasses GPT-4, highlighting the effectiveness of axiomatic training for causal reasoning on complex datasets.
@@
Overall, our work provides a new paradigm of teaching models causal reasoning through symbolic demonstrations of axioms or rules, which we call axiomatic training. Such symbolic data can be cheaply generated for multiple axioms and added to the finetuning data of language models. More generally, our results using axiomatic data contribute to the literature on causal learning from passive data (Lampinen et al., 2023). Our code repository can be accessed at: https://github.com/AniketVashishtha/Causal_Axioms.
@@ 2 Related Work
LLMs for Knowledge-Driven Causal Reasoning: Recent developments in Large Language Models (LLMs) have highlighted their potential for knowledge-driven causal discovery. Unlike traditional methods which focus on statistical patterns or correlations, LLMs utilize knowledge acquired through their pretraining to reason about and identify causal structures based on metadata of variables (Kıcıman et al., 2023; Ban et al., 2023; Long et al., 2023; Willig et al., 2022; Vashishtha et al., 2025). However, possibility of memorization of existing benchmarks in the pretraining of these LLMs has been a major criticism. As a result, recent work (Zečević et al., ) argues that LLMs are not actually performing causal reasoning, but simply learning correlations about causal facts. In addition, there are critical failure modes of using LLMs for causal discovery due to hallucinations or not obeying the acyclic constraint when generating graph edges (Vashishtha et al., 2025). To evaluate causal reasoning capabilities of LLMs, (Jin et al., 2024) and (Jin et al., 2023) propose formal causal inference evaluation benchmarks to infer direct and indirect causal relationships, and highlight the failure of LLMs in performing accurate causal reasoning.
Impact of Positional Encoding on Generalization: Length generalization capabilities of transformers has been studied in the past to better understand their different failure modes across various settings (Hupkes et al., 2020; Zhang et al., 2023; Furrer et al., 2021). Previous work (Kazemnejad et al., 2023; Bhattamishra et al., 2020; Haviv et al., 2022; Shen et al., 2023) emphasizes the impact of positional encoding in length generalization capability of transformers. To understand how transformers can be optimized for learning through axiomatic training and generalizing to unseen larger causal structures, we also examine different types of positional encoding such as no positional encoding (PE), Learnable PEs (Radford et al., 2018) and Sinusoidal PEs (Vaswani et al., 2017).
@@ 3.1 Axioms of Causality: Transitivity
$(X\nrightarrow Y|Z)\implies{}$ $(A\nrightarrow Y|Z){}\vee(X\nrightarrow A|Z)$
$\text{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }\forall A\notin X\cup Z\cup Y$
which can be simplified using the contrapositive. We call the LHS as Premise and the RHS as Hypothesis.
$\exists A\notin X\cup Y\cup Z\>\>s.t.\text{\ \ \ \ }$ $\underbrace{(X\rightarrow A|Z)\wedge(A\rightarrow Y|Z)}_{P:\text{premise}}$ (1)
@@
Definition 3.3 (Definition 1.2.3 in Pearl (2009)).
@@ 4.1 Training Data: Diversity is key
We train the model on data from simple causal graphs such as sequential chains with 3-6 nodes and evaluate its performance on more complex graphs (refer fig. 1). To enhance generalization, we introduce structured perturbations in the training data across three axes: node names, causal structure types, and the number of nodes in the causal graph.
@@
      Sequential: All causal edges are directed forward, forming a typical chain DAG, e.g., X $\rightarrow$ Y $\rightarrow$ Z.
@@ 4.2 Tokenization, Training Loss & Architecture
We train a decoder-based 67 million parameter model based on GPT-2’s architecture. The model has 12 attention layers, 8 attention heads and 512 embedding dimensions. We train the transformer model from scratch to ensure that the model has not seen such axioms in the pertaining step and thus requires a true correct understanding of axioms to perform well. Later we also tested on a pre-trained model which was fine-tuned on our dataset.
@@
Positional Encoding. In addition to tokenization and loss function, recent work (Kazemnejad et al., 2023) shows that the choice of positional encoding is important for generalizing a transformer to longer or complex inputs. Therefore, we evaluate different positional encodings on their impact on generalization in causal tasks: learnable (LPE) (Radford et al., 2018), sinusoidal (SPE) (Vaswani et al., 2017), rotary (RoPE) position encodings (Su et al., 2024), and no positional encoding (NoPE) (Kazemnejad et al., 2023; Haviv et al., 2022). See Appendix F for details.
Finetuning. Apart from training a transformer from scratch, we fine-tune a pre-trained language model (Llama-3-8b-Instruct (gra, 2024)) on our axiomatic training data. Finetuning a pretrained model like Llama, allows us to evaluate causal reasoning in a wider variety of contexts, leveraging its expansive vocabulary and the versatility provided by pretraining. While a model trained from scratch on axiomatic data is limited by its narrow vocabulary and structure, finetuning a pretrained model allows direct evaluation on benchmarks like Corr2Cause, which implicitly require the application of the target axioms, thus enabling us to stress test our framework in more diverse settings. Refer J for more details about the finetuning setup.
4.3 Evaluation setup: Assessing Axiomatic Learning
@@
    Node Name Shift: Testing the model’s performance on longer node names, from 1-3 characters in the training set to 8-10 characters, motivated by Jin et al. (2024) who show how change in node names can lead to generalization failure on causal tasks for language models.
@@
    Branching Factor: One of the most complex evaluation setups, with DAGs containing multiple branches, colliders, and forks. We define the branching factor as the ratio of the number of edges to the number of nodes. Thus, the branching factor for the training set is $\leq 1$. Then, we create a different evaluation set with multiple densely branched networks constructed using the Erdös-Rényi model, with different branching factors.
4.4 Benchmark evaluation
To test whether such simple axiomatic training is helpful in more complex scenarios, we evaluate our models on causal reasoning benchmarks, Corr2Cause (Jin et al., 2024) and CLEAR (Chen et al., 2024). Unlike past methods as in Jin et al. (2024), we do not finetune on these datasets directly. Instead, we finetune a language model on synthetic axiomatic data and evaluate whether the model generalizes to these external benchmarks. We discuss the results on these benchmarks in Section 7.
4.5 Baselines Using Existing LLMs
Given recent work on how LLMs can be leveraged for causal reasoning (Kıcıman et al., 2023; Vashishtha et al., 2025; Ban et al., 2023), we include language models such as GPT-4 (gpt-4-32k) (OpenAI, 2023), Gemini (gemini-pro) (Google DeepMind & Google Research, 2023) and Phi-3 (Phi-3-mini-128k-instruct) (Abdin et al., 2024) as baselines. Each model is significantly larger than our model and known to perform well on reasoning tasks, with the smallest model Phi-3 having 3.8B parameters (Li et al., 2023). Refer App. B and C.1 for zero and multi-shot prompts used. On evaluating the baselines for the transitivity task, we find that GPT-4 is significantly better. Hence, for the more complex d-separation task, we evaluate only GPT-4.
@@ 5.1 Training and Evaluation Datasets
Training Datasets. The training data consists of sequential chains of lengths from [3,6]. In addition to sequential chains, random flipping of edges is done with 0.5 probability. See Appendix G for details on these hyperparameters. Our training data consists of 175k axiom demonstrations. We use three versions of training data to evaluate the impact of different data perturbations.
@@
    Training Setup 1 (TS1): This setup comprises of 73k examples where the underlying graphs have random flipping and the remaining 101k causal graphs are sequential chains. Since flipping is done randomly across all consecutive pairs of nodes in the given chain, some complete reversals are also formed. In this training set, around 12k completely reversed chains are present.
@@
    Training Setup 2 (TS2): This setup comprises of more simple sequential chains (132k), while we decrease chains with random flipping (42k) to keep the overall size around 175k to see the effect of adding examples with complicated graphs on model’s performance.
Evaluation Datasets. In addition to the evaluation sets described earlier in Sec 4.3 (length generalization, node name shift, order of chains, and branching), we add another evaluation set that is a combination of three shifts.
$\mathbf{MultiEval_{SLR}}$ (Shuffling + Random Flipping + Length Sequence): This setup merges three kinds of complexities: shuffling of sentence for representing the sequences, each sequence having random flipping, and some sequences having longer length than sequences in training set (upto 9).
5.2 Results: Generalizing to Complex Causal Graphs